OpenStudy (puppylover52):

The formula d = rt gives the distance d traveled in time t at rate r. A bicyclist rides at a constant rate of 22 kmph. How many kilometers will he travel in 2.5 h?

OpenStudy (puppylover52):

plz help

OpenStudy (puppylover52):

asap

Vocaloid (vocaloid):

r = 22 t = 2.5 r*t = ?

OpenStudy (puppylover52):

um 22*2.5=550.0

OpenStudy (puppylover52):

is that correct

Vocaloid (vocaloid):

check your math again

Vocaloid (vocaloid):

22*2.5 = ?

OpenStudy (puppylover52):

i know i did

Vocaloid (vocaloid):

your answer is wrong. use a calculator if you are unsure.

Vocaloid (vocaloid):

what is 22 times 2.5?

OpenStudy (puppylover52):

55

Vocaloid (vocaloid):

good, so that's your answer

OpenStudy (puppylover52):

same thing

Vocaloid (vocaloid):

550.0 is not the same thing as 55...

Vocaloid (vocaloid):

that's like saying 100 is the same thing as 10...

OpenStudy (puppylover52):

yea if theres a decimal

Vocaloid (vocaloid):

550.0 is 550, not 55...

OpenStudy (puppylover52):

i was already gonna take the 0's away